OpenStudy (anonymous):

What is 27x2 − 13x2 A. 13 B. 13x2 C. 14 D. 14x2

OpenStudy (johnweldon1993):

since these both have x^2 ...they can be subtracted...however you don't touch the x^2 here you simply do 27 - 13 = ...? and whatever that equals...you put x^2 after it

OpenStudy (anonymous):

54-26=28=14*2

OpenStudy (anonymous):

14

OpenStudy (johnweldon1993):

right...27 - 13...is 14 but don't forget to bring down the x^2 with it so you would have 14x^2

OpenStudy (anonymous):

D

OpenStudy (anonymous):

think you both

OpenStudy (johnweldon1993):

no problem!